OpenStudy (anonymous):

Help me develope a 3*3 system of ODE IVP from this word problem: Suppose three chemical species are in a process of reaction, where every minute 20% of chemical A becomes chemical B, and 25% of chemical B become chemical C. There is no supply of chemical A. The reaction stops at chemical C. At present, there are 100 grams of chemical A, 20 grams of chemical B, and no chemical C. We are interested in the first 10 minutes of the reaction.

OpenStudy (blacksteel):

I'm not an expert at ODEs, but I think you'd get\[da(t)/dt = 0.8*a(t)\]\[db(t)/dt = 0.2*a(t)+0.75*b(t)\]\[dc(t)/dt = 0.25*b(t)\]

OpenStudy (anonymous):

so we wouldnt have c(t) in one of the equations?